Circular No. 7514

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMETS C/2000 U1, C/2000 U2, C/2000 U3, C/2000 U4 (SOHO) Further to IAUC 7508, D. Biesecker and D. Hammer report the dis-

coveries of four additional Kreutz sungrazing comets. All four were visible in the LASCO C2 coronagraph, and C/2000 U2 and C/2000 U3 were also seen in C3 images. C/2000 U1 and C/2000 U4 were found by M. Oates; C/2000 U2 was found by M. Boschat; and C/2000 U3 was found by J. Danaher. C/2000 U1 was described as faint, with no tail evident. C/2000 U2 was fairly bright, with a long, thin tail. The measurements were made by Hammer, and the reduced positions and orbits by B. G. Marsden are given on MPEC 2000-U38.

Comet  2000 UT  α_{2000}  δ_{2000}
C/2000 U1  Oct. 21.393  13^{h}37.^{m}0  −11^{o}40′
C/2000 U2  22.279  13 28.6  −13 13
C/2000 U3  24.029  13 41.4  −14 06
C/2000 U4  25.076  13 52.6  −13 09

SUPERNOVA 2000dx IN UGC 1775 M. Schwartz, Cottage Grove, OR; and W. D. Li, A. V. Filippenko, M.

Modjaz, and R. R. Treffers, University of California at Berkeley, report the discovery of a supernova during a collaboration of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126) using the 0.8-m Katzman Automatic Imaging Telescope (KAIT) and the Tenagra Observatory Supernova Search using the 0.8-m Tenagra II and 0.5-m Tenagra III automated telescopes (hereafter the Lick Observatory and Tenagra Observatory Supernova Searches, or LOTOSS). SN 2000dx was discovered and confirmed with the 0.5-m Tenagra III telescope on unfiltered images taken on Oct. 30.3 (mag ∼ 17.3) and 31.2 UT (mag ∼ 17.1). The new object was also confirmed on an unfiltered KAIT image taken on Oct. 31.3 (mag ∼ 17.1). SN 2000dx is located at α = 2^{h}18^{m}26^{s}.88, δ = +5^{o}39′02′′.0 (equinox 2000.0), which is 7′′.9 east and 12′′.3 south of the nucleus of UGC 1775. A KAIT image taken on Oct. 22.3 showed nothing at the position of SN 2000dx (limiting magnitude ∼ 18.5). Filippenko and Li add that inspection of CCD spectra (range 330–1000 nm) obtained on Nov. 1 with the Shane 3-m reflector at Lick Observatory reveals that SN 2000dx is of type Ia, probably a few days before maximum brightness.

2000 November 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green